14cm^3, each of solutions A and B with pH of 5 and 3 respectively are mixed. Determine the pH of the solution
1
Expert's answer
2019-10-21T07:17:18-0400

For solution "A" ,"[H^+](M_1)=10^{-5} mol\/L"

Volume"(V_1) =14\ cm^3=14 \times 10^{-3} dm^3=14 ml"

For solution B,"[H^+]=10^{-3} mol\/L"

Volume"(V_2)=14 cm^3=14 \ ml"

After mixing,Volume will be"(V_3)=14+14=28 ml"

Using molarity equation of mixing,

"M_1V_1+M_2V_2=M_3V_3"

"10^{-5}\times 14+10^{-3}\times 14=M_3\times 28"

"M_3=\frac{10^{-3}+10^{-5}}{2}=\frac{101}{200000}" mol/L

pH of solution"=-log(M_3)=3.30"